OpenStudy (anonymous):

Find the slope of the line passing through the points (5,3) and (2,2)?

OpenStudy (anonymous):

again IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

first find slope, then substitute into point slope formula

OpenStudy (anonymous):

slope = y2-y1 / x2-x1